A sum of $3248 is to be divided between two people in the ratio of 5 to 3. How much does each person receive?
I tried
3248* .60 = 1847 one person 1401 other person
I divided 5/3 to get ration % if 60 and other percent is 40 to create 100%
Thank you for your help.

Answer by gencore(1)

Let the first person with ratio 5 be A and the other with 3 be B. Add the ratio:5+3=8 the amount gotten by A will be 5/8 x $3248=$2030 and B will get 3/8 x $3248=$1218 the amount gotten by A and B is $2030 and $1218 respectively.